# Unsupervised Discovery of Long-Term Spatiotemporal Periodic Workflows in Human Activities

### v1

@@ 1 Introduction
To bridge this gap, we introduce a pioneering benchmark for unsupervised detection of long-term periodic spatiotemporal workflows in human activities. Specifically, we collected $580$ human activity sequences, with a small size but a rich diversity to cover most real-world periodic activities, such as factory production, exercise training, and shuttle routes. Our dataset covers rich modalities, including 3D body/hand pose sequences, 3D indoor trajectories, and 2D outdoor trajectories (Fig. 2). With our provided spatiotemporal features, we have designed three evaluation tasks. In the first task, we introduce an activity sequence consisting of more than three periods, requiring unsupervised identification of period boundaries and counting. The second task focuses on completion tracking, which involves using the extracted workflow to estimate the remaining portion of a new period when only part of that period is available. The final task utilizes the extracted workflow to localize procedural anomalies within a new period. These tasks correspond to real-world applications, such as assisting production line workers in optimizing schedules and reducing errors.
@@ 3 Our Benchmark
Evaluation Metrics. For task 1, we adopt Mean Absolute Percentage Error (MAPE) to evaluate period counting accuracy. This approach aligns with previous evaluation protocols in S-PAD [21, 75, 58, 56, 48], explicitly using its correct name. Specifically, suppose we have $N$ sequences, the estimated and ground truth (GT) period counts are denoted by $\{\hat{m}_{1},\hat{m}_{2},\cdots,\hat{m}_{N}\}$ and $\{m_{1},m_{2},\cdots,m_{N},\forall m_{i}\geq 3\}$, respectively. The MAPE is defined as:
@@ 4 Our Baseline Method
Although human activity occurs continuously in spatiotemporal space, its features can be discretized into tokenized representations. Ideally, activity tokens might be represented as textual semantic tags [104, 65]; however, despite advances in vision foundation models [3, 82, 100, 8], it is challenging to fully capture nuanced spatial-temporal relationships, intricate coordination, and procedural complexities that transcend verbal descriptions [46]. For instance, text alone cannot precisely convey the dynamic motions of two hands. To effectively manage activity tokens without additional model training, an alternative approach is to represent them using a finite set of symbols [32, 46, 33], with each symbol corresponding to a specific activity spatiotemporal state. Consequently, a transcript of human activity can be tokenized into a symbolic transcript (i.e., a string). Given a human activity transcript of length $T$ with frame-wise spatiotemporal features $\{\mathbf{x}_{t}\in\mathbb{R}^{n\times 1}\}_{t=0}^{T-1}$, we perform spatiotemporal clustering to organize these features into $K$ clusters. While $K$ can be automatically determined using non-parametric clustering methods [68], we simply use K-means clustering [39] for convenience. Empirically, a setting $K$ between $6$ and $14$ is sufficient to represent most workflows aimed at our task completion tracking and anomaly detection.
@@ 5 Experiments
Traditional unsupervised sequential mining methods struggle on our benchmark. Unlike previous datasets used for periodicity and anomaly detection studies, our dataset features low-contrast periodic patterns and complex workflows that significantly challenge existing approaches (Tab. 3(a)). For instance, in period counting, the MAPE scores of unsupervised sequential periodic detection methods we evaluated [25, 64, 36, 12] are over $50$ times higher than that of our baseline method. Similarly, in anomaly detection, the representative methods we assessed [29, 86, 88] perform nearly at the level of random guessing. Therefore, our benchmark presents a novel challenge to the research community, encouraging the development of new solutions.

### v2

@@ 1 Introduction
To bridge this gap, we introduce a pioneering benchmark for unsupervised detection of long-term periodic spatiotemporal workflows in human activities. Specifically, we collected 580580 human activity sequences, with a small size but a rich diversity to cover most real-world periodic activities, such as factory production, exercise training, and shuttle routes. Our dataset covers rich modalities, including 3D body/hand pose sequences, 3D indoor trajectories, and 2D outdoor trajectories (Fig. 2). With our provided spatiotemporal features, we have designed three evaluation tasks. In the first task, we introduce an activity sequence consisting of more than three periods, requiring unsupervised identification of period boundaries and counting. The second task focuses on completion tracking, which involves using the extracted workflow to estimate the remaining portion of a new period when only part of that period is available. The final task utilizes the extracted workflow to localize procedural anomalies within a new period. These tasks correspond to real-world applications, such as assisting production line workers in optimizing schedules and reducing errors.
@@ 3 Our Benchmark
Evaluation Metrics. For task 1, we adopt Mean Absolute Percentage Error (MAPE) to evaluate period counting accuracy. This approach aligns with previous evaluation protocols in S-PAD [21, 75, 58, 56, 48], explicitly using its correct name. Specifically, suppose we have NN sequences, the estimated and ground truth (GT) period counts are denoted by $\{\hat{m}_{1},\hat{m}_{2},\cdots,\hat{m}_{N}\}$ and $\{m_{1},m_{2},\cdots,m_{N},\forall m_{i}\geq 3\}$, respectively. The MAPE is defined as:
@@ 4 Our Baseline Method
Although human activity occurs continuously in spatiotemporal space, its features can be discretized into tokenized representations. Ideally, activity tokens might be represented as textual semantic tags [104, 65]; however, despite advances in vision foundation models [3, 82, 100, 8], it is challenging to fully capture nuanced spatial-temporal relationships, intricate coordination, and procedural complexities that transcend verbal descriptions [46]. For instance, text alone cannot precisely convey the dynamic motions of two hands. To effectively manage activity tokens without additional model training, an alternative approach is to represent them using a finite set of symbols [32, 46, 33], with each symbol corresponding to a specific activity spatiotemporal state. Consequently, a transcript of human activity can be tokenized into a symbolic transcript (i.e., a string). Given a human activity transcript of length $T$ with frame-wise spatiotemporal features $\{\mathbf{x}_{t}\in\mathbb{R}^{n\times 1}\}_{t=0}^{T-1}$, we perform spatiotemporal clustering to organize these features into $K$ clusters. While $K$ can be automatically determined using non-parametric clustering methods [68], we simply use K-means clustering [39] for convenience. Empirically, a setting $K$ between 66 and 1414 is sufficient to represent most workflows aimed at our task completion tracking and anomaly detection.
@@ 5 Experiments
Traditional unsupervised sequential mining methods struggle on our benchmark. Unlike previous datasets used for periodicity and anomaly detection studies, our dataset features low-contrast periodic patterns and complex workflows that significantly challenge existing approaches (Tab. 3(a)). For instance, in period counting, the MAPE scores of unsupervised sequential periodic detection methods we evaluated [25, 64, 36, 12] are over 5050 times higher than that of our baseline method. Similarly, in anomaly detection, the representative methods we assessed [29, 86, 88] perform nearly at the level of random guessing. Therefore, our benchmark presents a novel challenge to the research community, encouraging the development of new solutions.